For the following molecule with a coefficient: 3Fe2(C204)3
Indicate how many atoms of FE?
Indicate how many atoms of C?
Indicate how many atoms of O?
1
Expert's answer
2020-11-10T14:01:49-0500

Fe "\to" no. of molecules(3) × no. of atoms in a molecule(2) = 6 atoms.


C "\to" no. of molecules(3) × no. of atoms in a molecule(2 × 3) = 18 atoms.


O "\to" no. of molecules(3) × no. of atoms in a molecule(4 × 3) = 36 atoms.